What current strength in amperes will be required to liberate 30g of iodine from potassium iodide solution in one hour?
 

Grade:12th pass

1 Answers

Arun
4 years ago
 

Applying E = Z × 96500 (E for Iodine = 126) 
126 = Z × 96500 
or Z = 126/96500 g 
Now, applying the formula

W = Z × I × t

Where W = 30 g, Z= 126/96500 t = 60×60 =3600 second 

I = 30x96500/126x3600= 6.382 Amp